OpenStudy (anonymous):

Determine the mass of the cube of aluminum (Al) that mesures 6cm on each side and has a density of 216g/cm^3 Show work plz

OpenStudy (anonymous):

Alright so do you know what the formula for density is?

OpenStudy (anonymous):

hello.....

OpenStudy (anonymous):

The formula for density is d = density m = mass v = volume sorry i crashed

OpenStudy (anonymous):

s'all good, i crashed a while ago. We need more RAM dude.

OpenStudy (anonymous):

\[D\frac{ M }{ V }\]

OpenStudy (anonymous):

remember the equal sign..

OpenStudy (anonymous):

XD ops

OpenStudy (anonymous):

So you got density, you don't have volume and you don't have mass. Can we find volume?

OpenStudy (anonymous):

find the mass of Al and multiply it w/ density?

OpenStudy (anonymous):

No

OpenStudy (anonymous):

What is the significance of 6cm in the question?

OpenStudy (anonymous):

on each side of the cube... the size

OpenStudy (anonymous):

Eh ill just tell you. Use the lengths of the cube to get volume. Hence, 6cm*6cm*6cm=V

OpenStudy (anonymous):

XD lol sorry im out of it today.

OpenStudy (anonymous):

Well did you get your answer?

OpenStudy (anonymous):

\[216=\frac{ ? }{ 216 }\]

OpenStudy (anonymous):

What is your volume? What is your density?

OpenStudy (anonymous):

D=216 V=216 (6*6*6=216)

OpenStudy (anonymous):

so your mass is?

OpenStudy (anonymous):

idk

OpenStudy (anonymous):

Rearrange the formula for mass

OpenStudy (anonymous):

M=D*V

OpenStudy (anonymous):

so your answer is?

OpenStudy (anonymous):

M=46656

OpenStudy (anonymous):

convert that to scientific notation and you're all good.

OpenStudy (anonymous):

4.6656*10^4

OpenStudy (anonymous):

make it 4.67*10^4 g good job

OpenStudy (anonymous):

Tnx ya rock :)